OpenStudy (anonymous):

You are painting a ceiling, part of which is covered by a round light fixture 2 feet in a diameter. The ceiling is 12 feet by 8 feet. find the area of the region you will paint

OpenStudy (anonymous):

how do u do it?

OpenStudy (anonymous):

area of the ceiling - the area of the round light

OpenStudy (anonymous):

So u just subtract and thats it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh ok thank:)

OpenStudy (anonymous):

any cat treat?

OpenStudy (anonymous):

Wait there are three numbers which is which

OpenStudy (anonymous):

area of the ceiling = 12*8

OpenStudy (anonymous):

area of the light = pi*r^2

OpenStudy (anonymous):

and the radius is 1 so -> pi*1^2=3.14